OpenStudy (anonymous):

8x-y+5z from the sum of 2x+3y - 6z and 5x-3y

OpenStudy (anonymous):

13331

OpenStudy (anonymous):

pls complete solution :)

OpenStudy (anonymous):

6971

OpenStudy (callisto):

the sum of 2x+3y - 6z and 5x-3y = (2x + 3y -6z) + (5x - 3y) = 2x + 5x +3y - 3y -6z = 7x - 6z To get 8x-y+5z from the sum of 2x+3y - 6z and 5x-3y, we need to add something. Let that something be A [ (2x + 3y -6z) + (5x - 3y) ] + A = 8x-y+5z ( 7x - 6z ) + A = 8x - y + 5z Subtract both sides by 7x ( 7x - 6z ) + A - 7x = 8x - y + 5z -7x A - 6z = x - y + 5z Add 6z to both sides A - 6z + 6z = x -y + 5z + 6z A = x - y + 11z So, to get 8x-y+5z from the sum of 2x+3y - 6z and 5x-3y, we need to add x - y + 11z.

OpenStudy (callisto):

I hope I didn't misinterpret the question :(

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you did fine

OpenStudy (anonymous):

THANKS

OpenStudy (anonymous):

so whats the answer..?

OpenStudy (anonymous):

15

OpenStudy (callisto):

@tishara429 Can you explain your answer? I don't understand how you got your answer.

OpenStudy (anonymous):

add

OpenStudy (callisto):

Do you mind explaining it detailedly? Sorry that I'm slow!

OpenStudy (anonymous):

ok you add the number which equals the answer u got and then you will have the correct answer so thats how you solve the question.

OpenStudy (callisto):

But there are some unknowns there, you can't add them directly because they are NOT the like terms, right?

OpenStudy (anonymous):

oh i see then u have to divide